OpenStudy (anonymous):

Solve the following in [0,2pi) tan2(theta)cos(theta)=cos(theta)

OpenStudy (anonymous):

I attempted to do this problem but im not really sure if i did it right i will post what i did so far i really hope you can help me

OpenStudy (raden):

sorry, i change theta be "x" it make easier :) tan2xcosx=cosx tan2xcosx-cosx=0 cosx(tan2x-1)=0 for zeros, cosx=0 or tan2x-1=0 * cosx=0 --->x=pi/2, 3pi/2 ** tan2x-1=0 tan2x=1/2 (hmmm... not special degree for tan has value equal 1/2)

OpenStudy (anonymous):

ok for cos would i write it theta=pi/2+2pi k and theta=3pi/2+2pi k?

OpenStudy (anonymous):

and for tan2x-1=0 can you do it with tan2x=1 instead of zero please what would that be?

OpenStudy (anonymous):

hello?

OpenStudy (raden):

sorry, my conection very slow :( yeaa, i mistaken just knew (above)... tan2x-1=0 tan2x=1 tan2x=tanpi/4 tan2x=tan(k*pi+pi/4) 2x=(k*pi+pi/4 x=k*pi/2 + pi/8 tan2x=tan5pi/4 tan2x=tan(k*pi+5pi/4) 2x=k*pi+5pi/4 x=k*pi/2+5pi/8

OpenStudy (anonymous):

thank you so would pi/2,5pi/8,3pi/2,pi/8 be the only ones in the solution range [0,2pi)

OpenStudy (raden):

welcome :)

OpenStudy (anonymous):

are those that i written above correct ?

OpenStudy (anonymous):

i also have another question that i really need help on and if you can help me i would really appreciate it

OpenStudy (raden):

for x=k*pi/2 + pi/8 if k=0, --> x=pi/8 k=1, --> x=pi/2 + pi/8 = 5pi/8 k=2, --> x=pi + pi/8 = 9pi/8 k=3, --> x=3pi/2 + pi/8 = 13pi/8 k=4,5,6,so on not satisfied in interval [0,2pi)

OpenStudy (anonymous):

thank you can you help me with this next problem im really stuck on how to solve it

OpenStudy (raden):

yeps...

OpenStudy (anonymous):

ill write it in a new problem so then you can get another medal

OpenStudy (raden):

ok